OpenStudy (mrs.ambrose614):

@satellite73 can you help me

OpenStudy (templeguy):

what is it

OpenStudy (mrs.ambrose614):

Question 5 (Multiple Choice Worth 2 points) (05.05) The networking organization you joined is throwing a party. You are in charge of buying the chips, which cost $2.50 per bag and salsa, which costs $4 per jar. The chips & salsa budget you are given totals $60. The inequality 2.5x + 4y less than or equal to 60 represents the possible combinations of chips (x) and salsa (y) you can buy. graph of the inequality 2.5 x plus 4 y is less than or equal to 60 Which of the following does not represent a solution to the inequality? 10 bags of chips and 2 jars of salsa 20 bags of chips and 2 jars of salsa 14 bags of chips and 5 jars of salsa 18 bags of chips and 6 jars of salsa

OpenStudy (mrs.ambrose614):

Can u help m???

satellite73 (satellite73):

you need the graph of \(2.5x+4y\leq 60\)?

OpenStudy (mrs.ambrose614):

I already did that

satellite73 (satellite73):

oh ok that was the hard part, rest is easier

satellite73 (satellite73):

10 bags of chips and 2 jars of salsa\[2.5\times 10+4\times 2\leq 60\] is either true or false

OpenStudy (mrs.ambrose614):

So the AWNSER is??

satellite73 (satellite73):

i get \[25+8\leq 60\\ 33\leq 60\]which is true, so that one is possible

OpenStudy (mrs.ambrose614):

I'm being timed can you tell me the anwser I've been waiting all night

satellite73 (satellite73):

i cannot know without checking each one

satellite73 (satellite73):

it is probably this one 18 bags of chips and 6 jars of salsa

OpenStudy (mrs.ambrose614):

Do u know for sure

satellite73 (satellite73):

\[2.5\times 18+4\times6=?\]

OpenStudy (mrs.ambrose614):

60????

satellite73 (satellite73):

i get 69, which is bigger than 60 so that one is your answer

OpenStudy (mrs.ambrose614):

Question 6 (Multiple Choice Worth 1 points) (05.05) Which of the following inequalities matches the graph? graph of an inequality with a dashed line through the points (0, 3) and (1, 9) with shading above the line −6x + y < 3 6x + y < 3 6x − y < −3 The correct inequality is not listed Can you help me with this